Biosphere 2 Research Wiki » xe.blog.code.title

xe.blog.code.title

Last modified by Michael Sibayan on 2015/02/13 12:22
Comments (0) · Attachments (0) · History

Mar 24 2020

Corrected center bay PE102 leak data

PE102 Center bay for the leak was wired incorrectly, hence, it was never recording properly since 2018 till Jan 2020.

Since this PE102 only measured the leak from the seepage, it only exhibited low flow rate and was never above the minimum flowrate of the PE102.

However, due to a recent accidental discovery of the seepage leak in the center bay with the -2 and -1 seepage locations clogged, it had increased the flowrate of the leak to a point that was greater than the Novalynx tipping bucket measurable flowrate and within the measurable flowrate of the PE102.

The Database has been updated with the corrected PE102 center bay leak flowrate.

Feb 12 2020

3130 Load cell temperature

We have started logging the 3130 load cell temperature in order to potentially perform temperature compensation for the thermal drifts.

East - 2/2/2020 ~3:50pm

Center & West - 2/11/2020 ~3:15pm

Feb 06 2020

LEO Center Seepage Leak position testing

We recently noticed the flowrate increase with the LEO center seepage leak position that was higher than the novalynx could handled and had found that -2 and -1 seepage drains from the slop were clogged.  

The following were completed on 2/6/2020 to see how the clogging affected the seepage leak flowrate.

1st test with -2 plugged from 9:50am to 10:33am

2nd test with -1 plugged from 1:35pm to 2:20pm

Jan 31 2020

Center Seepage Clog

Unclogged the following locations on LEO Center

1/28/2019 ->  -2

1/31/2019 -> -1

Jan 22 2020

PE102 and Novalynx maintenance.

West slope.

PE102:

  • -4 - funnel was not directly under the drain. Repositioned.
  • -2, 2 & 4 may need to be flushed/cleaned
  • Reseated the questionable PE102 connectors

Novalynx:

  • -2 - tipping bucket was stuck. Adjusted the screws holding the shaft so it moves freely.   

Center slope:
Novalynx:

  • -1 and leak - replaced bad proxy switches
  • 2 - tipping bucket was caught on a screw that was along the side of the body.   

.

Sep 26 2019

Tipping bucket calibration

Tipping bucket calibration testing on Sept 25th,  between 12:46 pm to 2:47 pm.  

Sep 25 2019

Out of the schedule rain on miniLEO

According to the schedule, rain on miniLEO stopped at 16:30 on Sept 24th. Rain started again at 17:35 unexpectedly and continued overnight until 8:05 in the morning. Rain scheduler in use was without PID.

Rainfall distribution testing on minLEO surface

Rain fall distribution was tested on miniLEO surface on September 22nd, 2019 between 9.52am to 10.07am and 10.57am  to 11.12 am. Sprinkler heads were adjusted after first test. Second test showed reasonably uniform distribution.

Sep 16 2019

1 sec/min Data logging Performance Issue

There was a performance issue with the main dsc.b2 server with generating log files for 1 minute and 1 second log files.

Affected time range: 9/11/2019 8:10 PM to 9/16/2019 10:40AM

Aug 14 2019

LEO CNR4 Center Bay West Sensor Re-purposed

The west most CNR4 Sensor (LEO-C_10_-2_0_CNR4) was removed from the Center Bay mast on Thursday August 8th between 8am and 10 am by Aaron Bugaj.

This sensor is temporary being used in the West bay for a remote sensing experiment.

Tags:
Created by Michael Sibayan on 2015/02/13 12:22

This wiki is licensed under a Creative Commons 2.0 license
XWiki Enterprise 6.0.1 - Documentation